OpenStudy (anonymous):

find the equation for the tangent line to f(x)= 3^2x+1 at a=-2

OpenStudy (gorv):

find df/dx

OpenStudy (anonymous):

what is that

OpenStudy (phi):

are you studying calculus ?

OpenStudy (anonymous):

yes but this pellet is hard as hell

OpenStudy (anonymous):

pellet

OpenStudy (phi):

can you use the equation editor to put in your equation? I can't tell what is the exponent

OpenStudy (phi):

is it \[ f(x)= 3^{2x+1} \]?

OpenStudy (anonymous):

\[f(x)= 3^{2x+1} at a=-2\]

OpenStudy (phi):

when x is in the exponent , think "e" in other words, we can write \[ 3= e^{\ln3} \] and your problem is \[ f(x) = \left( e^{\ln3}\right)^{2x+1} \\ f(x) = e^{ (2\ln3)x + \ln3} \]

OpenStudy (anonymous):

yea bro i have no clue what this means

OpenStudy (phi):

You can try Khan's site http://www.khanacademy.org/math/calculus/differential-calculus/derivative_intro/v/calculus--derivatives-1--new-hd-version he has lots of stuff on calculus. But if you don't know this, you are not going to learn it by me typing in this box. It will take a lot of work on your part.

OpenStudy (gorv):

just take log both side logf(x)= log(3)^(2x+1) logf(x)= (2x+1)*log3=2xlog3+log3 now differentiate with respect to x on differentiating 1/f(x) * f ' (x)=2*log3 f ' (x)=f(x)*2*log3 but f(x)= 3^2x+1 f ' (x)=3^(2x+1)*2*log3 at x=-2 f ' (-2)=3^(2*(-2)+1)*2*log3=3^(-3)*2*log3=(2log3)/27 this is slope of tangent

OpenStudy (gorv):

f(x)= 3^2x+1 at x=-2 f(-2)=y=3^(-4+1)=1/27 so it is passing through (-2,1/27) and slope=(2log3)/27